A person walks 25.00 north of east for 3.10 km. How far would she have to walk due north and due east to arrive at the same location?
1
Expert's answer
2020-10-13T07:09:13-0400

let the distance due east = "x"

"Cos25^{o}= \frac{x}{3.1}"

"x=3.1 \times Cos25^{o}= 2.81 Km"

let the distance due north "=y"

"Sin25^{o} = \frac{y}{3.1}"

"y= 3.1 \times Sin25^{o} = 1.31 Km"